Covington Corporation purchased a vibratory finishing machine for $23,000 in year 0. The useful life of the machine is 10 years, at the end of which the machine is estimated to have a salvage value of zero. The machine generates net annual revenues of $6,000. The annual operating and maintenance expenses are estimated to be $500. If​ Covington's MARR is 14​%,

how many years will it take before this machine becomes​ profitable? Assume that the cash flows occur continuously throughout the year.
Answer #1

Find out when the NPV = 0

Net revenue = 6000-500 =5500

NPV = -23000+5500*(P/A,14%,n)

0 = -23000+5500*(P/A,14%,n)

23000/5500 = P/A,14%,n

4.1818 = P/A,14%,N

N = 6.7 years or 7 years